OpenStudy (anonymous):

The speed of the current in a river is 6 mph. A ferry operator who works that part of the river is looking to buy a new boat for his business. Every day, his route takes him 22.5 miles against the current and back to his dock, and he needs to make this trip in a total of 9 hours. He has a boat in mind, but he can only test it on a lake where there is no current. How fast must the boat go on the lake in order for it to serve the ferry operator’s needs? @amistre64 @Mertsj @e.mccormick @zepdrix @agent0smith help i dont understand?

OpenStudy (mertsj):

The problem is asking you to find the speed of the boat in still water. Do you understand that part?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

Let r = the speed of the boat in still water. When the boat is going with the current, its speed is r+6 When the boat is going against the current, its speed is r-6

OpenStudy (mertsj):

Let's say it takes him t hours to go the 22.5 miles with the current. Then , since rt = d, we have t(r+6)=22.5

OpenStudy (mertsj):

Then to go the 22.5 miles against the current will take him the rest of the 9 hours which is 9-t and so we have: (r-6)(9-t)=22.5

OpenStudy (mertsj):

Solve that system.

OpenStudy (anonymous):

so you want me to solve (r-6)(9-t)=22.5?